) A computer is infected with the Sasser virus. Assume that it infects 20 other computers within 5 minutes; and that these PCs and servers each infect 20 more machines within another five minutes, etc. How long until 100 million computers are infected?
Answer by fractalier(2101)

The computer viral growth is an exponential function. Call it N(t), where N is the number of infected computers and t is the time in minutes. Thus,
N(t) = 20 ^ (t/5)
Now set this equal to 100,000,000 and solve for t, using logs...
20 ^ (t/5) = 100000000
log [20 ^ (t/5)] = log (100000000)
(t/5)log (20) = 8
t = 40 / log 20 or about 30.74 minutes